OpenStudy (anonymous):

4x²y times 3x²y³ divided by 2xy

OpenStudy (anonymous):

i may be wrong, but i reckon it would come to 6x^3y^3 you have 4x^2y multiplied by 3x^2y^3 which = 3 times 4 times x^2 times x^2 times Y times y^3, the indices/powers add so this = 12x^4y^4 divide by 2xy. to begin with 12/2 =6 then the indices/powers subtract because you are dividing so x^4/x=x^3 and y^4/y = y^3 result = 6x^3y^3

OpenStudy (anonymous):

Thanks for the reply :) what does ^ stand for?

OpenStudy (anonymous):

to the power of, so x^2 means x squared and y^3 means y cubed :)